OpenStudy (anonymous):

Find the limit of the function algebraically: 3sin(4x) / sin(3x)

OpenStudy (lgbasallote):

approaching what?

OpenStudy (anonymous):

Oh shoot sorry, approaching 0.

OpenStudy (helder_edwin):

\[ \large \lim_{x\to0}\frac{3\sin(4x)}{\sin(3x)}= \lim_{x\to0}\frac{\frac{4\sin(4x)}{4x}}{\frac{\sin(3x)}{3x}} \]

OpenStudy (helder_edwin):

remember this \[ \large \lim_{t\to0}\frac{\sin t}{t}=1 \]

OpenStudy (anonymous):

Wow thats alot to take in. Hopefully Ill make sense of it. Thanks!

OpenStudy (helder_edwin):

u r welcome